two springs have force constants kand k2. Both are stretched till their elastic energies are equal .If the stretching forces are F1 and F2 then F1: Fis

Grade:10

1 Answers

Vikas TU
7 years ago
elastic energies are equal hence,
0.5k1x1^2 = 0.5k2x2^2
k1/k2 = (x1/x2)^2............................(1)
 
F1 = k1x1
and
F2 = k2x2
F1/F2 = k1x1/k2x2 => (k1/k2)^(3/2) {from eqn. (1)}
F1:F2 =>    k1^3/2  :   k2^3/2